# 2022-deformable-elastic-objects-UAVs

### v1

@@ I Introduction and Related Work
Deformable object manipulation is an important recent development in aerial robotics with potential applications ranging from fire fighting [1], and in general the manipulation of fluid conduits [2], to waterway maintenance [3], e.g., in case of oil-spill events [4]. Yet, for the challenges it involves [5], deformable object manipulation is still regarded as an open problem. An extensive survey on modeling, perception, and control techniques for deformable object manipulation is in [6]. This work especially focuses on the manipulation of cables through Uncrewed Aerial Vehicles (UAVs).
Aerial robotic manipulation has traditionally considered rigid cables or linear elastic cables, typically as tools to manipulate attached payloads. More recently, increasingly general models have been used to describe cable flexibility, too. This work considers a model that describes flexible and extensible cables, thus enlarging the application range. In the following, we provide a quick overview of related works.
@@ I-3 Flexible inextensible cable models
They allow cable bending but not length variations. [14] proposes a controller to stabilize a quadrotor with a hanging bendable cable; similarly, an adaptive control to account for unknown mass is proposed in [15]. [16] considers cooperative aerial manipulation through flexible inextensible cables, and [17] shows the differential flatness of a point-mass or a rigid body suspended below multiple quadrotors through flexible inextensible cables. Moreover, the differential flatness of single- and multi-robot systems connected to these cables is shown in [1]. A different modeling approach than the lumped-parameter ones seen so far is adopted in [18] and [2], where two-quadrotors are connected by a cable modeled as a catenary curve. Such a model describes the kinematics of the cable through a small number of parameters but is mostly limited to static conditions.
@@ I-4 Flexible and extensible cable models
They allow the cable to bend and change its length. While the use of such models is found in manipulation with ground robots [19, 6, 20], it is still mostly unexplored in the aerial robotic manipulation domain. [3] demonstrated the differential flatness of the discretized cable model composed of springs interconnected by lumped-mass passive spherical joints and attached between two quadrotors. The model incorporates external viscous forces, too, and it is the same used in this work. Other methods consider a distributed-parameter cable model, such as [21], which applies back-stepping control to suppress the vibrations of a hose transported by a fixed-wing vehicle. To control the cable shape, [22] applies proper orthogonal decomposition to the partial differential equations that describe the dynamics of the cable, combined with the ordinary differential equations that describe the dynamics of the attached quadrotor.
@@ II Model
We consider deformable elastic cables and model them through a series of $n$ point masses interconnected by springs through passive spherical joints. Let us define =$\mathcal{I}=\{1,\cdots,n\}$ the set of the indexes that we use to refer to the point masses. Hence, we indicate the position of mass $m_{i}$ expressed in $\mathcal{F}_{W}$ as $\bm{p}_{i}\in\mathbb{R}^{3}$, for $i\in\mathcal{I}$. For class $(a)$, we also define as $\bm{p}_{0}=\bm{0}$ the position of the end of the cable attached to the ground; without loss of generality, attached to $O_{W}$. As said, the masses are connected to elastic links modeled as linear springs; such a modeling approach for deformable objects has proven effective when dealing with small deformations [6]. For the spring that has one end attached to the point in position $\bm{p}_{j}$ and the other to the one in position $\bm{p}_{j+1}$ with $j\in\mathcal{I}\setminus\{n\}$ we define a stiffness coefficient $k_{j}$ and rest length ${l}_{0j}$ f. The force exerted by the $j^{\rm{th}}$ elastic segment on the $j^{\rm{th}}$ point mass is indicated as $\bm{f}_{j}.$ Consequently, $-\bm{f}_{j}$ is the force exerted by the same segment on the $(j+1)^{\rm{th}}$ mass. We assume that $\bm{f}_{j}\neq\bm{0}$. For systems in class $(a)$, we also define a spring with elastic coefficient $k_{0}\in\mathbb{R}^{+}$ and rest length ${l}_{00}\in\mathbb{R}^{+}$ that connects the first mass to the ground and exerts a force $\bm{f}_{0}$ on it. Considering a simplified Hook’s law to model the springs, we have that $\forall i\in\mathcal{I}\setminus\{n\}$,
$\bm{f}_{i}=-$ ${k}_{i}\left[\left(\bm{p}_{i}-\bm{p}_{i+1}\right)-{l}_{0i}\frac{\bm{p}_{i}-\bm{p}_{i+1}}{||\bm{p}_{i}-\bm{p}_{i+1}||},\right]$ (1)
where we indicated with ||||||*||| | ∗ | | the 2-norm. Force $\bm{f}_{i}$ acts along the unit vector ${\frac{\bm{p}_{i}-\bm{p}_{i+1}}{||\bm{p}_{i}-\bm{p}_{i+1}||}=\frac{\bm{f}_{i}}{||\bm{f}_{i}||}.}$ Moreover, for systems $(a)$ $\bm{f}_{0}={k}_{0}\left[\bm{p}_{1}+{l}_{00}\frac{\bm{p}_{1}}{||\bm{p}_{1}||}\right]$, where $-\bm{f}_{0}$ acts on the mass $m_{1}$.
@@
For $j\in\mathcal{R}$ and $i\in\mathcal{I}\setminus\mathcal{R}$, and defining $\bm{f}_{0}=\bm{0}$ for classes $(b),(c)$, the dynamics equations of systems $(a)-(c)$ are as follows
@@ III Differential Flatness
Figure 2: Simulation results for a system of class $(a1)$ in (a) and class $(b)$ in (b). In the animation screenshot, the desired output trajectory and the output points are in green. In the plots, $x-,y-,z-$ coordinates of the output error.
@@
    $\bar{\bm{f}}_{1}=-{k}_{1}\left[\left(\bm{p}_{1}-\bm{p}_{2}\right)-{l}_{01}\frac{\bm{p}_{1}-\bm{p}_{2}}{||\bm{p}_{1}-\bm{p}_{2}||}.\right]$ (8)
    We note that the unit vector expressing the direction of the force is $\frac{\bm{p}_{1}-\bm{p}_{2}}{||\bm{p}_{1}-\bm{p}_{2}||}=\frac{\bar{\bm{f}}_{1}}{||\bar{\bm{f}}_{1}||}$. Then, we rewrite (8) as
    $\bm{p}_{2}=\bm{p}_{1}+\frac{\bar{\bm{f}}_{1}}{{k}_{1}}-{l}_{01}\frac{\bar{\bm{f}}_{1}}{||\bar{\bm{f}}_{1}||},$ (9)
@@ IV Simulation Results
Figure 5: $||\bm{p}_{i}^{meas}-\bm{p}_{i}^{sim}||$[m] Top: measured (red) and estimated (blue) cable trajectories subject to the same quadrotor trajectories for the narrow fast Test A (left) and the fast Test B (right). Below are the corresponding output error plots, where $\bm{p}_{i}^{sim}$ is the estimated value of the output $\bm{p}_{i}$ and $\bm{p}_{i}^{meas}$ the measured one. ||.||||.||| | . | | is the 2-norm
@@ V-A Cable Parameter Identification
In the experiment, a 1-meter cable with mass $7\cdot 10^{-3}$ kg has been uniformly discretized into 6 points, including the two extremes where the robots would be attached. Hence, $n=6$ and $n_{R}=2$. Reflective markers have been attached to each point mass to record their position using a motion capture system; the data was recorded at 100 Hz. No quadrotors were used for the collection of the parameter identification dataset. While that may result in some unmodeled components such as aerodynamic interference of the propellers on the cable, it allows for a wider range of motion of the cables without worrying about instability issues and avoids the need for planning complex robot trajectories, which may result in being unfeasible. Instead, the cable was manually perturbed via the extremities. As the variability of the dataset affects the parameter identification, a wide range of configurations and velocities have been generated. Validation with the robots at the extremities of the cable is shown in the last part of this section.
@@
$\begin{split}&\underset{\bm{\theta},\hat{\mathbf{x}}_{i}\hat{\mathbf{x}}_{i_{c}}}{\text{min}}\sum_{i=2}^{n-1}\left(\sum_{t=0}^{t_{\text{end}}}\frac{1}{\lambda}\left(\hat{\mathbf{x}}_{i}(t)-\mathbf{x}_{i}(t)\right)^{\text{T}}\mathbf{W}\left(\hat{\mathbf{x}}_{i}(t)-\mathbf{x}_{i}(t)\right)+\right.\\ &\left.+\frac{1}{1-\lambda}\left(\hat{\mathbf{x}}_{i_{c}}(t)-\hat{\mathbf{x}}_{i}(t)\right)^{\text{T}}\mathbf{W}\left(\hat{\mathbf{x}}_{i_{c}}(t)-\hat{\mathbf{x}}_{i}(t)\right)\right)\end{split}$ (12)
@@
[…]
@@
$\overline{{\mathbf{e}}_{3}}$ 0.443 m 71.1 mm 0.367 m 0.498 m
$\overline{\mathbf{e}_{4}}$ 0.384 m 0.572 m 0.292 m 0.409 m
@@
$\overline{\mathbf{e}_{3}}$ 0.289 m 0.274 m
$\overline{\mathbf{e}_{4}}$ 0.145 m 0.209 m
TABLE II: $\overline{\bm{e}_{i}}$ is the average of $||\bm{p}^{sim}_{i}-\bm{p}^{meas}_{i}||_{2}$, $i={3,4}$.
@@ V-B Cable Manipulation Control
In this section, we assess the performance of a cable-control strategy, based on the discrete cable model, closing the control loop on the cable outputs. We test a simple controller that integrates into the computation of the UAV trajectories, based on the differential flatness as explained in Section III, an integral term on the output error. That is similar to what we proposed in [3] for systems $(c1)$ and tested solely through simulations so far. Specifically, given as the two output the positions of two consecutive points on the discrete cable model, $\bm{p}_{i}$ and $\bm{p}_{i+1}$, and given their desired trajectories as well, we compute the corresponding desired force $\bm{f}_{i}^{d}$ and we arrive at computing $\bar{\bm{f}}_{i+1}^{d}$. We define $\bm{e}_{i}=\bm{p}_{i}^{d}-\bm{p}_{i}$ and write the analogous of (9) as ${\bm{p}_{i+2}^{d}=\bm{p}_{i+1}^{d}+\frac{\bar{\bm{f}}_{i+1}}{{k}_{i+1}}-{l}_{0i+1}\frac{\bar{\bm{f}}_{i+1}^{d}}{||\bar{\bm{f}}_{i+1}^{d}||}+K^{I}\int_{0}^{t}\bm{e}_{i}(\tau)\rm{d}\tau,}$ where $\tau$ is an auxiliary variable, and $K^{I}\in\mathbb{R}^{3\times 3}$ is a positive-definite diagonal matrix with elements to the diagonal set to $0.2$ in the tests. The same procedure is applied when computing $\bm{p}_{i-1}$ and so on, proceeding along the other side of the cable. Note that, with this method, the UAV references are continuously recomputed online based on output feedback. The controller runs offboard and the references are sent at 100 Hz via radio communication to the quadrotors. The use of a different cable and different output point locations induces parameter uncertainties. In this first closed-loop test, a $3.6{\rm g}$ weight, $1.55{\rm m}$ long cable is iterated as the controlled object, while the two output points are located in the middle of the cable with $0.22{\rm m}$ distance. In the second closed-loop test, a $2.6{\rm g}$ weight, $1.24{\rm m}$ long cable is implemented, while the two output points are shifted toward one side of the cable. The results of two closed-loop experiments are in Fig. 6, in which the same reference output as in Test B (narrow) was used. The experimental results as well as the simulations can be found in the video attachment.
@@ VI Discussion
The proposed model showed an average error on each coordinate of the output lower than $0.02\rm{m}$ in the dataset used for parameter identification. During validation test A, the wide configurations outperformed the narrow ones in both execution velocities. Intuitively, an increased stretch in the cable reduces its dangling motion. The results also show that increasing the velocity of the execution increases the error. Overall, the model performance was satisfactory in the validation tests A and B, with the highest norm of the average output error registered in the narrow and fast test A and equal to $0.071\rm{m}$ for $\bm{p}_{3}$ and $0.57\rm{m}$ for $\bm{p}_{4}.$ The sensor (the motion capture system in this case) would be imputable to part of the error; however, that can be neglected here as its accuracy is a few millimeters. Note that all available data were used for the parameter estimation with no division between training and validation datasets. This is justified by the low order of the discrete model, chosen a priori for the sake of efficiency. However, high-order models, possibly requiring different data-driven estimation methods, would require careful handling of the training dataset to test against overfitting.
The closed-loop experiments showed that the controller is able to reduce the output error effectively. In the first closed-loop experiment (Fig.6a), before the controller was activated, the average tracking error was $0.373\rm{m}$ and $0.347\rm{m}$ for $\bm{p}_{3}$ and $\bm{p}_{4}$, respectively. After the controller was switched on, these two values decreased to $0.108\rm{m}$ and $0.100\rm{m}$. Similarly, for the second closed-loop experiment (Fig. 6b), the tracking errors for $\bm{p}_{3}$ and $\bm{p}_{4}$ were $0.216\rm{m}$ and $0.324\rm{m}$ in the open loop and decreased to $0.088\rm{m}$ and $0.095\rm{m}$ with the closed loop.

### v2

# Manipulation of Elasto-Flexible Cables with Single or Multiple UAVs

@@ I Introduction and Related Work
Deformable object manipulation is an important recent development in aerial robotics with potential applications ranging from fire fighting [1], and in general the manipulation of fluid conduits [2], to waterway maintenance, e.g., floating litter collection [3, gonzalezmorgado2025multirobotaerialsoft] or handling oil-spill events [4]. Yet, for the challenges it involves [5], deformable object manipulation is still regarded as an open problem. An extensive survey on modeling, perception, and control techniques for deformable object manipulation is in [6]. This work especially focuses on the manipulation of cables through Uncrewed Aerial Vehicles (UAVs).
Aerial robotic manipulation has traditionally considered rigid cables or linear elastic cables, typically as tools to manipulate attached payloads. More recently, increasingly general models have been used to describe cable flexibility, too. In the following, we provide a quick overview of related works.
@@ I-3 Flexible inextensible cable models
They allow cable bending but not length variations. [14] proposes a controller to stabilize a quadrotor with a hanging bendable cable; similarly, an adaptive control to account for unknown mass is proposed in [15]. [16] considers cooperative aerial manipulation through flexible inextensible cables, and [17] shows the differential flatness of a point-mass or a rigid body suspended below multiple quadrotors through flexible inextensible cables. Moreover, the differential flatness of single- and multi-robot systems connected to these cables is shown in [1]. A different modeling approach than is adopted in [18, 2, gonzalezmorgado2025multirobotaerialsoft], where two-quadrotors are connected by a cable modeled as a catenary curve or a parabola. Such models describe the kinematics of the cable through a small number of parameters but are mostly limited to quasi-static conditions.
@@ I-4 Flexible and extensible cable models
They allow the cable to bend and change its length, thus enlarging the application range. While the use of such models is found in manipulation with ground robots [19, 6, 20], it is still mostly unexplored in the aerial robotic manipulation domain. [3] demonstrated the differential flatness of the discretized cable model composed of springs interconnected by lumped-mass passive spherical joints and attached between two quadrotors. The model incorporates external viscous forces, too, and it is the same used in this work. Other methods consider a distributed-parameter cable model, such as [21], which applies back-stepping control to suppress the vibrations of a hose transported by a fixed-wing vehicle. To control the cable shape, [22] applies proper orthogonal decomposition to the partial differential equations that describe the dynamics of the cable, combined with the ordinary differential equations that describe the dynamics of the attached quadrotor.
@@ II Model
We consider deformable elastic cables and model them through a series of $n$ point masses interconnected by springs through passive spherical joints. Let us define $\mathcal{I}=\{1,\cdots,n\}$ the set of the indexes that we use to refer to the point masses. Hence, we indicate the position of mass $m_{i}$ expressed in $\mathcal{F}_{W}$ as $\bm{p}_{i}\in\mathbb{R}^{3}$, for $i\in\mathcal{I}$. For class $(a)$, we also define as $\bm{p}_{0}=\bm{0}$ the position of the end of the cable attached to the ground; without loss of generality, attached to $O_{W}$. As said, the masses are connected to elastic links modeled as linear springs; such a modeling approach for deformable objects has proven effective when dealing with small deformations [6]. For the spring that has one end attached to the point in position $\bm{p}_{j}$ and the other to the one in position $\bm{p}_{j+1}$ with $j\in\mathcal{I}\setminus\{n\}$ we define a stiffness coefficient $k_{j}\in\mathbb{R}^{+}$ and rest length ${l}_{0j}\in\mathbb{R}^{+}$. The force exerted by the $j^{\rm{th}}$ elastic segment on the $j^{\rm{th}}$ point mass is indicated as $\bm{f}_{j}.$ Consequently, $\bm{f}_{j}$ is the force exerted by the same segment on the $(j+1)^{\rm{th}}$ mass. We assume that $\bm{f}_{j}\neq\bm{0}$. For systems in class $(a)$, we also define a spring with elastic coefficient $k_{0}\in\mathbb{R}^{+}$ and rest length ${l}_{00}\in\mathbb{R}^{+}$ that connects the first mass to the ground and exerts a force $\bm{f}_{0}$ on it. Considering a simplified Hook’s law to model the springs, we have that $\forall i\in\mathcal{I}\setminus\{n\}$,
[…]
@@
For $j\in\mathcal{R}$ and $i\in\mathcal{I}\setminus\mathcal{R}$, and defining $\bm{f}_{0}=\bm{0}$ for classes $(b),(c)$, the dynamics equations of systems (a)(c)(a)-(c) are as follows
@@ III Differential Flatness
Figure 2: Simulation results for a system of class $(a1)$ in (a) and class $(b)$ in (b). In the animation screenshot, the desired output trajectory and the output points are in green. In the plots, x,y,zx-,y-,z- coordinates of the output error.
@@
[…]
    which is now an equation in the only unknown $\bm{p}_{2}$. We retrieve $\bm{p}_{2}$ as a function of the flat outputs from (9), and its derivatives are a function of the flat output derivatives by differentiating (9).
@@ IV Simulation Results
Figure 5: $\bm{p}_{i}^{meas}-\bm{p}_{i}^{sim}||$[m] Top: measured (red) and estimated (blue) cable trajectories subject to the same quadrotor trajectories for the narrow fast Test A (left) and the fast Test B (right). Below are the corresponding output error plots, where $\bm{p}_{i}^{sim}$ is the estimated value of the output $\bm{p}_{i}$ and $\bm{p}_{i}^{meas}$ the measured one. $||.||$ is the 2-norm
@@ V-A Cable Parameter Identification
In the experiment, a 1-meter cable with mass $7\cdot 10^{-3}$ kg has been uniformly discretized into 6 points, including the two extremes where the robots would be attached. Hence, $n=6$ and $n_{R}=2$. Reflective tape has been attached to each point mass to record their position using a motion capture system; the data was recorded at 100 Hz. No quadrotors were used for the collection of the parameter identification dataset. While that may result in some unmodeled components such as aerodynamic interference of the propellers on the cable, it allows for a wider range of motion of the cables without worrying about instability issues and avoids the need for planning complex robot trajectories, which may result in being unfeasible. Instead, the cable was manually perturbed via the extremities. As the variability of the dataset affects the parameter identification, a wide range of configurations and velocities have been generated. Validation with the robots at the extremities of the cable is shown in the last part of this section.
@@
$\begin{split}&\underset{\bm{\theta},\hat{\mathbf{x}}_{i}\hat{\mathbf{x}}_{i_{c}}}{\text{min}}\sum_{i=2}^{n-1}\left(\sum_{t=0}^{t_{\text{end}}}\frac{1}{\lambda}\left(\hat{\mathbf{x}}_{i}(t)-\mathbf{x}_{i}(t)\right)^{\text{T}}\mathbf{W}\left(\hat{\mathbf{x}}_{i}(t)-\mathbf{x}_{i}(t)\right)+\right.\\ &\left.+\frac{1}{1-\lambda}\left(\hat{\mathbf{x}}_{i_{c}}(t)-{\mathbf{x}}_{i}(t)\right)^{\text{T}}\mathbf{W}\left(\hat{\mathbf{x}}_{i_{c}}(t)-{\mathbf{x}}_{i}(t)\right)\right)\end{split}$ (12)
@@
$m_{i}$ $1.16\cdot 10^{-3}\text{kg}$ ${l}_{01}$ $0.1950\,\text{m}$
${l}_{02}$ $0.1942\,\text{m}$ ${l}_{03}$ $0.1827\,\text{m}$
${l}_{04}$ $0.1943\,\text{m}$ ${l}_{05}$ $0.1977\,\text{m}$
$k_{1}$ $11.312\,\text{N/m}$ $k_{2}$ $5.411\,\text{N/m}$
$k_{3}$ $15.519\,\text{N/m}$ $k_{4}$ $7.008\,\text{N/m}$
$k_{5}$ $14.477\,\text{N/m}$ $c$ $0.002\,\text{N s/m}$
@@
$\overline{{\mathbf{e}}_{3}}$ 0.044 m 0.071 m 0.037 m 0.050 m
$\overline{\mathbf{e}_{4}}$ 0.038 m 0.057 m 0.029 m 0.041 m
@@
$\overline{\mathbf{e}_{3}}$ 0.030 m 0.027 m
$\overline{\mathbf{e}_{4}}$ 0.014 m 0.021 m
TABLE II: $\overline{\bm{e}_{i}}$ is the average of $\bm{p}^{sim}_{i}-\bm{p}^{meas}_{i}||_{2}$, i=3,4i={3,4}.
@@ V-B Cable Manipulation Control
[…]
@@ VI Discussion
The proposed model showed an average error on each coordinate of the output lower than $0.02\rm{m}$ in the dataset used for parameter identification. During validation test A, the wide configurations outperformed the narrow ones in both execution velocities. Intuitively, an increased stretch in the cable reduces its dangling motion. The results also show that increasing the velocity of the execution increases the error. Overall, the model performance was satisfactory in the validation tests A and B, with the highest norm of the average output error registered in the narrow and fast test A and equal to $0.071\rm{m}$ for $\bm{p}_{3}$ and $0.057\rm{m}$ for $\bm{p}_{4}.$ Part of the error would be attributable to the sensor (the motion capture system in this case); however, that can be neglected here as its accuracy is a few millimeters. Note that all available data were used for the parameter estimation with no division between training and validation datasets. This is justified by the low order of the discrete model, chosen a priori for the sake of efficiency. However, high-order models, possibly requiring different data-driven estimation methods, would require careful handling of the training dataset to test against overfitting.
The closed-loop experiments showed that the controller is able to reduce the output error effectively. In the experiment shown in Fig.6a, before the controller was activated, the average tracking error was $0.373\rm{m}$ and $0.347\rm{m}$ for $\bm{p}_{3}$ and $\bm{p}_{4}$, respectively. After the controller was switched on, the values decreased to $0.108\rm{m}$ and $0.100\rm{m}$. Similarly, for the experiment reported in Fig. 6b, the open-loop tracking errors for $\bm{p}_{3}$ and $\bm{p}_{4}$ were $0.216\rm{m}$ and $0.324\rm{m}$ and decreased to $0.088\rm{m}$ and $0.095\rm{m}$ when closing the control loop.